A penny is 1.55 mm thick.Imagine that you stack pennies on top of each other every day. Each day, you double the amount of pennies in the stack. This means that on day 0, there is 1 penny. On day 1, there are 2 pennies. On day 2, there are 4 pennies, etc.if this pattern continues, in how many days will the stack of pennies be about 1 meter tall?
Answer by jorel555(1290)

If n equals the number of days, then:
1.55*2^n=1000
2^n=645.16
2^9=512, 2^10=1024
Your stack will be over a meter tall after the 9th day!!!!!!!!!!